The number of complaints per day, X, received by a cable TV distributor has the probability distribution 2. 0 .4 .2 a) Find the expected number of complaints per day 0C6. 4)I(o.3)2 (0.1) 3( 0,2) b) Find the standard deviation of the number of complaints What is the approximate probability that the distributor receives a total of more than 100 complaints in 90 days? c)
Answer #1

(a) The computation table for the frequency distributi on 0 0.4 1 0.3 2 0.2 3 0.1 0.3 0.4 0.3 0.3 0.8 0.9 4 The mean of the distribution is E (X)-Ax 2x f(x) b) The formul a for calculating the variance is The standard deviation is Therefore, the standard deviation is σχ

c) The approximate probability that the distributor receives a total of more than 100 complaints in 90 days is 100 Since, on
Answer #2

SOLUTION :


a.


p = f(x) 


E(x) 

= ∑ f(x) * x  

= 0.4*0 + 0.3*1 + 0.2*2 + 0.1*3 

= 1.0  company / day (ANSWER).


b.


Variance 

= ∑ f(f)( x - E(x))^2

= 0.4(0 - 1)^2 + 0.3(1 - 1)^2 + 0.2(2 - 1)^2 + 0.1(3 - 1)^2

= 0.4 + 0 + 0.2 + 0.4

= 1.0 


So,


SD 

= sqrt(Variance) 

= sqrt(1.0)

= 1.0 company / day (ANSWER).


c.


Company per day = 100/90 = 10/9 


P(x > 10/9) 

= P(z > (10/9 - 1)/ 1)

= P(z > 0.1111)

From ND table 

= 1 - 0.5442

= 0.4558 (ANSWER).

answered by: Tulsiram Garg